OpenStudy (anonymous):

A bridge is built in the shape of a semielliptical arch. It has a span of 112 feet. The height of the arch 26 feet from the center is to be 12 feet. Find the height of the arch at its center.

OpenStudy (anonymous):

Might be able to solve this one using standard form of ellipse equation: \[\large (x-h)^2/a^2+(y-k)^2/b^2=1.\]

OpenStudy (anonymous):

which one is h and k and b and a I need some one to explain me that

OpenStudy (anonymous):

The way I'm thinking of it, if you put the center of the ellipse at (0,0), then h=k=0, so that simplifies things. If the span is 112 feet, then the semi-major axis in the x direction is 56, that would be 'a' I think you have to use the given information to solve for 'b.'

OpenStudy (anonymous):

and c is?

OpenStudy (anonymous):

b^2= a^2 -c^2

OpenStudy (anonymous):

I wouldn't worry about that. I don't think that information is needed. If you model the ellipse as \[\large \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] With a=56 feet, and the point on the ellipse, (26,12) as (x,y) then if you solve that equation for b, that should be the answer you are looking for. 

OpenStudy (anonymous):

okay give me a second yo finish it

OpenStudy (anonymous):

13.55 ft

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Oh, I haven't worked it out yet, gimme a sec to check it...